Hi please help with the following problem!
5 (x-2) ≥ 7x
Thank you so much

Found 2 solutions by fractalier, Boreal:
Answer by fractalier(6550)

5(x-2) >= 7x
5x - 10 >= 7x
-10 >= 2x
-5 >= x
or
x <= -5

Answer by Boreal(15235)

5 (x-2) ≥ 7x
5x-10 >=7x; subtract 7x from both sides
-2x-10>=0
-2x>=10
x<= -5
Check the answer
x=-6
5(-8)>=-42, that is true.
x=-4
5(-6)>=-28, that is not true.